please help me solve this equation. y-1=2/7x.
(?, 5-7) find x
the part that keeps messing me up is getting rid of the 2/7 from the x.
I am doing 5/7-1= 2/7x(7/2)
(7/2)5/7-1=x
5/2-1=x
getting x=3/2 but it doesn't make sense when I plug it back into the problem. what am I doing wrong?

Answer by Alan3354(69443)

please help me solve this equation. y-1=2/7x.
(?, 5-7) find x
the part that keeps messing me up is getting rid of the 2/7 from the x.
I am doing 5/7-1= 2/7x(7/2)
(7/2)5/7-1=x
5/2-1=x
getting x=3/2 but it doesn't make sense when I plug it back into the problem. what am I doing wrong?
============================
It's not clear what you want to do.
What is (?,5-7)?
-----------------
You have 1 eqn in 2 variables. You can solve for x or y in terms of the other variable.
=======================================
=======================================
please help me solve this equation. y-1=2/7x.
(?, 5-7) find x
the part that keeps messing me up is getting rid of the 2/7 from the x.
I am doing 5/7-1= 2/7x(7/2)
(7/2)5/7-1=x
5/2-1=x
getting x=3/2 but it doesn't make sense when I plug it back into the problem. what am I doing wrong?
============================
It's not clear what you want to do.
What is (?,5-7)?
-----------------
You have 1 eqn in 2 variables. You can solve for x or y in terms of the other variable.
================================
y-1=2/7x.
(x,5/7) find x
Multiply by 7 to eliminate fractions
7y-7 = 2x
Sub for y
5-7 = 2x
x = -1